(root node)

ccit(0)   iso(1)
org(3)
dod(6)
internet(1)

directory(1)   mgmnt(2)
mib-2(1)

system(1)
sysDescr(1)   sysObjectID(2)   sysContact(4)   sysName(5)   sysLocation(6)   sysServices(7)   sysORLastChange(8)   sysORTable(9)

interfaces(2)
ifTable(2)
ifEntry(1)
ifIndex(1)_{1   2   3   4   5}

ifDescr(2)_{1   2   3   4   5}

ifType(3)_{1   2   3   4   5}

ifMtu(4)_{1   2   3   4   5}

ifSpeed(5)_{1   2   3   4   5}

ifPhysAddress(6)   ifAdminStatus(7)    ifOperStatus(8)    ifLastChange(9)     ifInOctets(10)    ifInUcastPkts(11)   ifInNUcastPkts(12)   ifInDiscards(13)    ifInErrors(14)   ifInUnknownProtos(15)   ifOutOctets(16)   ifOutUcastPkts(17)   ifOutNUcastPkts(18)   ifOutDiscards(19)   ifOutErrors(20)     ifOutQLen(21)      ifSpecific(22)
1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5    1   2   3   4   5   1   2   3   4   5   1   2   3   4   5    1   2   3   4   5    1   2   3   4   5   1   2   3   4   5     1   2   3   4   5    1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5

at(3)   ip(4)   icmp(5)   tcp(6)   udp(7)   egp(8)   transmission(10)   snmp(11)

experimental(3)                                                                            private(4)
                                        enterprises(1)

                              ibm(2)   cisco(9)   hp()   ict(11400)

                                                      security(5)   snmpV2(6)

   joint(2)